OpenStudy (anonymous):

julie recently drove to visit her parents who love 72 miles away. on her way there her average speed was 14 miles per hour faster than on her way home. if julie spent a total of 3 hours driving, find the two rates.

OpenStudy (anonymous):

72 [miles]/ x[miles/h] => time she need to go home 72 [miles} / (x+14) [miles/h] => time to go to her parent's. add then =3h x is the variable

OpenStudy (anonymous):

\[72/x + 72/(x+14)=3\]

OpenStudy (anonymous):

same denominator \[72(x+14)+72x=3*x(x+14)\]\[72*(x+14+x)=3x*(x+14)\]\[72*(2x+14)=3x²+42x\]\[144x+1008=3x²+42x\]\[3x²-102x-1008=0\] using \[\Delta=b²-4ac\] 2 solutions \[x=-b-\sqrt(\Delta)/2a\] or \[x=-b+\sqrt(\Delta)/2a\] x=-8 or 42 (-8 impossible because it's a speed) so x=42mph and 56mph if we check the solution 72/42 +72/56 = (72*56+72*42)/(56*42)=3 Job done